OpenStudy (anonymous):

Part 1: Explain, in complete sentences, the effect of the difference of squares pattern on the multiplication of radicals. Part 2: Give an example.

OpenStudy (highschoolmom2010):

If, for instance, we have \[a+\sqrt{b}\] and we multiply it with its conjugate \[a-\sqrt{b}\] where a is a real number and b is a positive number which is not a perfect square, then all we need remember is the pattern of "difference of two squares", like so : \[(x+y)(x-y)=x^2-y^2\] In our case, x=a and \[y=\sqrt{b}\] So the result of multiplying them is \[a^2-(\sqrt{b})=a^2-b\] In other words, the effect of the "difference of two squares" pattern is the loss of the radical sign, typically used to simplify equations. For example \[(2+\sqrt{3})(2-\sqrt{3})=2^2-(\sqrt{3})^2=4-3=1\]